# GraphLoRA: Empowering LLMs Fine-Tuning via Graph Collaboration of MoE

### v1

@@ 4.2 Expert Distinction Strategy
The core idea of MoE is to utilize the unique capability of each expert to collaborate toward superior outcomes. Hence, we optimize the assigned weights by graph router to approximate the Poisson distribution. Our aim is to elicit the distinct capabilities of different experts. Specifically, for the output vector of the graph router, denoted as $\mathbf{o}_{r}\in\mathbb{R}^{N}$. Each dimension of $\mathbf{o}_{r}$ represents the assigned weight for each expert to deal with the input information. As the order of experts has no practical meaning, we sort values of each dimension in the vector $\mathbf{o}_{r}$ in descending order to obtain a new vector $\mathbf{v}_{r}\Leftrightarrow\textit{sort}(\mathbf{o}_{r})$. The distribution of $\mathbf{v}_{r}\in\mathbb{R}^{N}$ is optimized to approximate the Poisson distribution vector $\mathbf{v}_{\textit{poisson}(\lambda,i)}$. The Kullback–Leibler (KL) divergence distance is used to calculate the loss function, defined as:
$\mathbf{v}_{\textit{poisson}(\lambda,i)}=\frac{\lambda^{i}}{i!}e^{-\lambda},$ (6)
@@
where $i=\{1,2...,N\}$ and $N$ is the number of experts. $\mathbf{v}_{\textit{poisson}(\lambda,i)}$ is the vector follows the Poisson distribution and $\lambda$ is the learning parameter.
@@ 4.3 Load Balance Strategy
Apart from the distinct capabilities of different experts, one important factor in MoE training is keeping the load balance of all experts. Otherwise, greater routing weights on a small number of experts in the early stages of fine-tuning will result in a rapid localized optimization problem. In GraphLoRA, we propose the Normal distribution-based load balance strategy. Specifically, as only Top-K experts are activated in dealing with input information, we calculate the cumulative weights of each expert to represent the activation frequency distribution of all experts in the fine-tuning process. And normalized the cumulative weights to construct the activation frequency vector, denoted as $\mathbf{v}_{a}$ in the next feed-forward step. Our aim is to make the activation frequency vector $\mathbf{v}_{a}$ follow a natural normal distribution rather than the absolute equality in existing MoE literature (Li et al. 2024a; Zoph et al. 2022). Formally, the Normal distribution-based load balance loss function can be defined as:
$\mathbf{v}_{\textit{normal}(\mu,\sigma,i)}=\frac{1}{\sqrt{2\pi\sigma}}\exp(-\frac{{(i-u)}^{2}}{2\sigma^{2}}),$ (8)
@@
where $i=\{1,2...,N\}$ and $N$ is the number of experts. $\mathbf{v}_{\textit{normal}(\mu,\sigma,i)}$ follows the Normal distribution, $\mu=\frac{N}{2}$ is the mean of all samples, and $\sigma$ is learning parameter optimized in the fine-tuning process.
@@ 5.1 Experimental Settings
To evaluate the performance of our framework, we use four widely used public datasets, including the question-answering task, i.e., ARC-Challenge (Clark et al. 2018), OpenBookQA (Mihaylov et al. 2018), SIQA (Sap et al. 2019), and task classification task in BoolQ (Clark et al. 2019) dataset. These datasets cover the evaluations on different domains of LLMs, such as factual knowledge from Wikipedia, natural science, science facts, and social interactions.
  • ARC-Challenge. It is released by the Allen Institute for AI and is widely used to evaluate the reading comprehension and reasoning abilities of LLMs. The dataset consists of multiple-choice questions derived from U.S. elementary school science exams.
  • BoolQ. It is a question-answering dataset and contains Yes/No type answers extracted from real search queries. Each question is described by a relevant passage of text. This dataset evaluates the natural language understanding capability of LLMs, particularly in handling short-form question answering and textual reasoning.
  • OpenBookQA. It is designed to evaluate the ability of LLMs to understand elementary-level science knowledge. It consists of multiple-choice questions. Each question is accompanied by a small set of ”facts” that are required to answer the questions.
  • SIQA. Social Interaction QA is a benchmark for multiple-choice question answering focused on social intelligence reasoning. It consists of questions that require models to understand and reason about social situations, such as interpreting human behavior, emotions, and intentions. Each question contains a brief introduction with one correct answer in three answer choices.
Table 1: The comparisons of model Accuracy and Stability measured by standard deviation (Std). The smaller the standard deviation, the greater the stability of the model. The underline represents the SOTA baseline method. The best results on Accuracy and Stability are highlighted in bold. For the Stability evaluation, the best experimental results in sparse MoE methods (except for LoRAMoE) are marked with *.
LLMs Method ARC-C($\uparrow$) BoolQ($\uparrow$) OBQA($\uparrow$) SIQA($\uparrow$) ARC-C Std.($\downarrow$) BoolQ Std.($\downarrow$) OBQA Std.($\downarrow$) SIQA Std.($\downarrow$)
Llama3 LoRAMoE 76.77 74.48 87.33 79.66 0.25 0.66 1.01 0.39
MING-MoE 77.28 72.95 86.47 79.48 0.91 0.91 1.42 0.31
MoLA 76.74 73.50 84.00 78.70 0.47 0.34 1.44 0.90
MixLoRA 77.36 75.43 87.20 79.53 0.52 0.99 1.13 0.15
GraphLoRA 77.56 75.90 88.13 80.48 0.26 0.17 0.98 0.37
Qwen2 LoRAMoE 83.50 74.80 90.53 80.54 0.44 0.03 0.23 1.15
MING-MoE 83.99 74.21 91.00 80.67 1.51 0.62 2.00 1.12
MoLA 83.19 74.48 89.93 80.59 1.34 0.75 1.79 0.83
MixLoRA 84.41 74.77 90.00 80.31 0.69 0.90 0.69 1.90
GraphLoRA 85.10 75.40 91.67 80.98 0.48 0.57 0.64 0.24
Yi-1.5 LoRAMoE 84.33 72.89 91.73 80.94 3.09 0.38 0.42 0.36
MING-MoE 84.58 73.32 90.07 81.56 0.61 0.58 0.50 0.39
MoLA 84.47 72.26 90.07 81.25 0.61 0.21 1.03 0.31
MixLoRA 84.36 73.32 91.80 81.89 1.75 0.39 0.40 0.31
GraphLoRA 85.32 74.23 91.33 82.24 0.52 0.32 0.40 0.28
(a) Llama3-8B
(b) Qwen2-7B
(c) Yi1.5-9B
Figure 2: The comparisons of model Efficiency of different MoE methods. The x-axis is the average of model accuracy on four datasets, and the y-axis is the number of trainable parameters.
@@
To verify the effectiveness of GraphLoRA, we compare the performance of typical PEFT MoE methods on three different base LLMs. Three popular open-source LLMs, i.e., Llama3-8B [^1] https://github.com/meta-llama/llama3, Qwen2-7B (Yang et al. 2024a), and Yi-1.5-9B (Young et al. 2024) are used in our experiments.
  • LoRAMoE (Dou et al. 2023b). It is a representative dense MoE model. A localized balancing constraint is used to alleviate the knowledge-forgetting problem in the model updating process.
  • MING-MoE (Liao et al. 2024). It is a MoE model designed for medical multi-task learning. It is a typical sparse MoE architecture without any constant loss function.
  • MoLA (Gao et al. 2024). It is a recently proposed MoE-based PEFT model that applies different numbers of experts in router functions in different layers. Specifically, more experts are used at higher layers of the transformer block.
  • MixLoRA (Li et al. 2024a). It is the SOTA method of PEFT in MoE literature. It adopts the MoE-LoRA architecture in the FFN layer of LLMs. To address the imbalance load problem, it uses an average auxiliary load balance loss.
  • GraphLoRA. Different from MixLoRA, we propose a novel graph router in GraphLoRA. It takes advantage of the collaboration of all experts learned by graph neural networks. Two coordination strategies, i.e., the expert distinct strategy and load balance strategy, are proposed to enhance the capabilities of all experts.
@@
All methods are implemented in PyTorch. For each baseline method, a grid search is applied to find the optimal settings. These include learning rate from $\{0.1,0.01,0.001,0.0001,0.00001\}$, number of experts from $\{4,8,12,16\}$, rank of LoRA from $\{1,2,4,8,16\}$, Top-K experts from $\{1,2,3,4,5\}$. We report the result of each method with its optimal hyperparameter settings on the validation data. In our model, we adopt GCN as the graph neural network. The number of aggregation layers in GCN is 2, and the hidden dimension of GCN layers is 256. The edge density $\beta$ to construct the MoE graph is $0.1$. The $\alpha$ in Eq. 1 is 4 in the LoRA component. The coefficients of the Poisson distribution loss function (see in Eq. 7) and Normal distribution loss function (see in Eq. 9) are set to 0.005 and 8 in the final loss function in LLM fine-tuning. More details of experimental implementation are available at our code link.
5.2 Main Results
We compare the performance of different MoE methods from three aspects, i.e., Accuracy, Stability, and Efficiency, to make comprehensive evaluations. The results on four datasets with three different base LLMs are shown in Table 1 and Fig. 2. We repeat each experiment under five different random seeds, and report the mean value of the model accuracy. The standard deviation of each model is used to measure the stability of method. The efficiency is evaluated by the size of learning parameters in the fine-tuning process of LLMs.
Accuracy Evaluation.
Model performance on accuracy metric is shown in Table 1. We can see that: (1) GraphLoRA achieves the best model performance in most cases on four datasets with different base LLMs, showing the effectiveness of the collaborations of multiple experts learned on the MoE graph. (2) The sparse MoE architecture with an auxiliary load balance loss in MixLoRA performs better than both the sparse MoE and dense MoE methods in many cases, showing the importance of enhancing the coordination of experts. (3) The performance of the methods varies across different datasets. For example, for the easier task with higher model accuracy in the OpenBookQA dataset, the dense MoE method LoRAMoE performs better in Llama3 and Yi-1.5. This indicates that the localized optimization of sparse MoE architecture may reduce the model capability and may lose its advantages in easy task learning. With effective expert collaboration in GraphLoRA, our method consistently exhibits the best performance in various downstream tasks.
Stability Evaluation.
The biggest challenge in MoE fine-tuning research is the instability of model performance caused by the imbalance load problem. As shown in Table 1, the smaller the standard deviation, the greater the stability of the model. We can see that: (1) The standard deviation of GraphLoRA is the smallest in most cases compared with the sparse MoE methods. It achieves comparable stability as the LoRAMoE with a dense MoE architecture, showing the superiority of our framework in dealing with the imbalance load problem. (2) The stability of different MoE methods varies on different downstream datasets, and it is also influenced by the base LLMs. The dense MoE method is relatively stable but suffers the cost of memory and computational overhead due to all experts being activated in the learning process.
Efficiency Evaluation.
We calculate the number of learning parameters used in each method. The comparisons of their average performance on four datasets and parameter size in three base LLMs are shown in Fig. 2. We can see that our GraphLoRA achieves the best model performance with the least number of trainable parameters. This phenomenon is mainly due to the fact that the rank of LoRA required in our framework is much lower than that of other methods. The effective expert cooperation mechanism in our graph router reduces the number of parameters required by each expert, enabling our method to achieve both effectiveness and efficiency.
5.3 Experimental Analysis
We conduct ablation studies to show the effectiveness of the graph router and two coordination strategies used in GraphLoRA. We further analyze the effects of hyper-parameters that may influence the model performance.
@@ 5.3 Experimental Analysis
Figure 3: The model performance of degradation variants of GraphLoRA on Qwen2-7B.
(a) Expert number
(b) Top-K
(c) Rank of LoRA
(d) Edge density
ARC-Challenge
(e) Expert number
(f) Top-K
(g) Rank of LoRA
(h) Edge density
BoolQ
@@
Ablation Studies.
@@
The graph router in GraphLoRA allocates weight for each expert based on the collaboration information learned from the MoE graph. Besides, the Poisson distribution-based expert distinction strategy and normal distribution-based load balance strategy are adopted to empower the capability of each expert and their collaborations. To verify the effectiveness of each component, we remove the graph router, Poisson distinct loss (see in Eq. 7), and Normal balance loss (see in Eq. 9) respectively. The performance comparisons of the degradation variants (-Graph), (-Normal), (-Poisson) on four datasets are shown in Fig. 3. We can see that all three components make contributions to the model performance. The impact of removing the graph router is obvious in all datasets in both model accuracy and stability aspects, indicating the importance of using the MoE graph to enhance expert collaborations. Besides, the Poisson distribution loss plays a crucial role in the fine-tuning process, showing the necessity of maintaining the distinct capability of each expert. The normal distribution loss contributes to keeping the load balance of all experts and substantially improves the performance of GraphLoRA.
Hyper-Parameter Analysis.
@@
The performance of GraphLoRA is affected by many hyper-parameters. We conduct analysis experiments to show the effects of the important hyper-parameters, including the number of experts used in MoE, the number of Top-K experts, the rank in the LoRA component, and the edge density in MoE graph construction. We present the results of ARC-Challenge and BoolQ datasets on Qwen2-7B in Fig. 4. We have the following observations.
@@
  • The number of experts. The number of experts varies from $\{4,8,12,16,32\}$. We can see that our model achieves the best performance with 8 experts. The increase in the number of experts may potentially complicate the collaboration process of experts and enlarge the imbalance problem.
@@
  • Top-K experts. The number of experts activated by the router function with the Top-K largest assigned weights. The $K$ varies from $\{1,2,3,4,5\}$ in our experiments, we find our model performs the best with $K=2$.
@@
  • Rank of LoRA. The hyper-parameter rank $r$ in LoRA controls the number of parameters that are updated in the fine-tuning process. The rank $r$ varies from $\{1,2,4,8,16,32\}$. Our model with $r=2$ achieves the best model performance which is less than the rank $r=4$ and $r=8$ required in MixLoRA in two datasets.
@@
  • Edge density. The edge density controls the connection among expert nodes. The edge density is computed as the ratio of the number of edges in the MoE graph compared to the edges in the fully connected graph. The edge density $\beta$ varies from $\{5\%,10\%,20\%,50\%,80\%,100\%\}$. We find that with only $10\%$ connected edges, the collaboration information shared among experts facilitates our model to achieve the best results.

### v3

@@ 4.2 Expert Distinction Strategy
The key aspect of the MoE framework lies in leveraging the unique capabilities of each expert to collaborate effectively, thereby achieving enhanced overall performance. Hence, to elicit the distinct capabilities of different experts, we optimize the assigned weights by the graph router to approximate a Poisson distribution. Specifically, for the output vector of the graph router, denoted as $\mathbf{o}_{r}\in\mathbb{R}^{N}$. Each dimension of $\mathbf{o}_{r}$ represents the assigned weight for each expert to deal with the input information. As the order of experts has no practical meaning, we sort values of each dimension in the vector $\mathbf{o}_{r}$ in descending order to obtain a new vector $\mathbf{v}_{r}\Leftrightarrow\textit{sort}(\mathbf{o}_{r})$. The distribution of $\mathbf{v}_{r}\in\mathbb{R}^{N}$ is optimized to approximate the Poisson distribution vector $\mathbf{v}_{\textit{poisson}(\lambda,i)}$. The Kullback–Leibler (KL) divergence distance is used to calculate the loss function, defined as:
$\mathbf{v}_{\textit{poisson}(\lambda,i)}=L_{norm}(\frac{\lambda^{i}}{i!}e^{-\lambda}),$ (6)
@@
where $i=\{1,2...,N\}$ and $N$ is the number of experts. $\mathbf{v}_{\textit{poisson}(\lambda,i)}$ is Poisson distribution vector with L1-normalization operation (i.e., $L_{norm}$) and $\lambda$ is the learning parameter.
@@ 4.3 Load Balance Strategy
Apart from the distinct capabilities of different experts, one important factor in MoE training is keeping the load balance of all experts. Otherwise, greater routing weights on a small number of experts in the early stages of the fine-tuning process will result in a rapid localized optimization problem. In GMoE, we propose the Normal distribution-based load balance strategy. Specifically, as only Top-K experts are activated in dealing with input information, we calculate the cumulative weights of each expert to represent its activation frequency. Then construct the activation frequency vector of all experts by normalizing the cumulative weights, denoted as $\mathbf{v}_{a}$ in the next feed-forward step. Our aim is to make the activation frequency vector $\mathbf{v}_{a}$ follow a natural normal distribution rather than the absolute equality in existing MoE literature (Li et al., 2024a; Zoph et al., 2022). Formally, the Normal distribution-based load balance loss function can be defined as:
$\mathbf{v}_{\textit{normal}(\mu,\sigma,i)}=L_{norm}(\frac{1}{\sqrt{2\pi\sigma}}\exp(-\frac{{(i-u)}^{2}}{2\sigma^{2}})),$ (8)
@@
where $i=\{1,2...,N\}$ and $N$ is the number of experts. $\mathbf{v}_{\textit{normal}(\mu,\sigma,i)}$ is the normalized vector that follows the Normal distribution, $\mu=\frac{N}{2}$ is the mean of all samples, and $\sigma$ is learning parameter optimized in the fine-tuning process.
@@ 5.1 Experimental Settings
To evaluate the performance of our framework, we use four representative public datasets, including the question-answering task, i.e., ARC-Challenge (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), SIQA (Sap et al., 2019), and task classification task in BoolQ (Clark et al., 2019) dataset. These datasets cover the evaluations on different domains of LLMs, such as factual knowledge from Wikipedia, natural science, science facts, and social interactions.
Table 1: The comparisons of model Accuracy and Stability measured by standard deviation (Std). The smaller the standard deviation, the greater the stability of the model. The underline represents the SOTA baseline method. The best results on Accuracy and Stability are highlighted in bold. For Stability evaluation, the superior outcomes of our approach compared with sparse MoE methods (excluding LoRAMoE) are emphasized in italics.
[…]
LLMs Method ARC-C BoolQ OBQA SIQA Avg. ARC-C BoolQ OBQA SIQA Avg.
Llama3 LoRAMoE 76.77 74.48 87.33 79.66 79.56 0.25 0.66 1.01 0.39 0.58
MING-MoE 77.28 72.95 86.47 79.48 79.05 0.91 0.91 1.42 0.31 0.89
MoLA 76.74 73.50 84.00 78.70 78.24 0.47 0.34 1.44 0.90 0.79
MixLoRA 77.36 75.43 87.20 79.53 79.88 0.52 0.99 1.13 0.15 0.70
GMoE 77.56 75.90 88.13 80.48 80.52 0.26 0.17 0.98 0.37 0.45
Qwen2 LoRAMoE 83.50 74.80 90.53 80.54 82.34 0.44 0.03 0.23 1.15 0.46
MING-MoE 83.99 74.21 91.00 80.67 82.47 1.51 0.62 2.00 1.12 1.31
MoLA 83.19 74.48 89.93 80.59 82.05 1.34 0.75 1.79 0.83 1.18
MixLoRA 84.41 74.77 90.00 80.31 82.37 0.69 0.90 0.69 1.90 1.04
GMoE 85.10 75.40 91.67 80.98 83.29 0.48 0.57 0.64 0.24 0.48
Yi-1.5 LoRAMoE 84.33 72.89 91.73 80.94 82.47 3.09 0.38 0.42 0.36 1.06
MING-MoE 84.58 73.32 90.07 81.56 82.38 0.61 0.58 0.50 0.39 0.52
MoLA 84.47 72.26 90.07 81.25 82.01 0.61 0.21 1.03 0.31 0.54
MixLoRA 84.36 73.32 91.80 81.89 82.84 1.75 0.39 0.40 0.31 0.71
GMoE 85.32 74.23 91.33 82.24 83.28 0.52 0.32 0.40 0.28 0.38
@@
To verify the effectiveness of GMoE, we compare the performance of typical PEFT MoE methods on three popular open-source LLMs, i.e., Llama3-8B [^1] https://github.com/meta-llama/llama3, Qwen2-7B (Yang et al., 2024), and Yi-1.5-9B (Young et al., 2024).
LoRAMoE (Dou et al., 2023b).
It is a representative dense MoE model. A localized balancing constraint is used to alleviate the knowledge-forgetting problem in the model updating process.
MING-MoE (Liao et al., 2024).
It is a typical sparse MoE architecture without constant loss functions and designed for medical multi-task learning.
MoLA (Gao et al., 2024).
It is a recently proposed MoE-based PEFT model that applies different numbers of experts in router functions in different layers. More experts are used at higher layers of the transformer block.
MixLoRA (Li et al., 2024a).
It is the SOTA method of PEFT in MoE literature. It adopts the MoE-LoRA architecture in the FFN layer of LLMs. To address the imbalance load problem, it uses an average auxiliary load balance loss.
GMoE.
Different from MixLoRA, we propose a novel graph router. It takes advantage of the collaboration of all experts learned by graph neural networks. Besides, two coordination strategies, i.e., the expert distinct strategy and load balance strategy, are proposed to enhance the capabilities of all experts.
@@
All methods are implemented in PyTorch with an A800-80G. For each baseline method, a grid search is applied to find the optimal settings. These include learning rate from $\{0.1,0.01,0.001,0.0001,0.00001\}$, number of experts from $\{4,8,12,16\}$, rank of LoRA from $\{1,2,4,8,16\}$, Top-K experts from $\{1,2,3,4,5\}$. We report the results of each method with its optimal hyperparameter settings on the validation data. In our model, we adopt GCN as the graph neural network (GNN). The number of aggregation layers in GCN is 2, and the hidden dimension of GCN layers is 256. The edge density $\beta$ to construct the MoE graph is $0.1$. The $\alpha$ in Eq. 1 is 4 in the LoRA component. The coefficients of the Poisson distribution loss function and Normal distribution loss function are set to 0.005 and 8 in the final loss function in LLM fine-tuning.
@@
Figure 2: The model performance of degradation variants of GMoE on Qwen2-7B.
Expert number
Top-K
Rank of LoRA
Edge density
Initial edge
(a) ARC-Challenge Dataset
Expert number
Top-K
Rank of LoRA
Edge density (%)
Initial edge
(b) BoolQ Dataset
@@
5.2 Main Results
We make comprehensive evaluations of different MoE methods from both Accuracy and Stability evaluations. We repeat each experiment under five different random seeds, and report the mean value of accuracy. The standard deviation of each model is used to measure the stability of different models. As shown in Table 1. The smaller the standard deviation, the greater the stability of the model. We can see that:
(1) The sparse MoE architecture with an auxiliary load balance loss in MixLoRA performs better than both the sparse MoE (i.e., MING-MoE, MoLA) and dense MoE methods (i.e., LoRAMoE) in many cases, showing the importance of enhancing the coordination of experts.
(2) The performance of the methods varies across different datasets. For example, for the easier task with higher model accuracy in the OpenBookQA dataset, the dense MoE method (i.e., LoRAMoE) performs better in Llama3 and Yi-1.5. This indicates that localized optimization of sparse MoE may reduce the model’s capability and lose its advantages in easy task learning.
(3) GMoE achieves the best performance in most cases on four datasets, showing the effectiveness of the collaborations of multiple experts learned on the MoE graph.
(4) Among sparse MoE methods (MING-MoE, MoLA, MixLoRA), GMoE exhibits the smallest standard deviation in most cases, achieving stability comparable to LoRAMoE, which is a dense MoE activating all experts for balance but with high computational costs.
(a) Llama3-8B
(b) Qwen2-7B
(c) Llama3-8B
(d) Qwen2-7B
Figure 4: Efficiency comparisons across MoE methods. Illustrated with Trainable parameters and Throughput (samples/second) on the x-axis and average accuracy across four datasets on the y-axis.
5.3 Ablation Studies
@@
We perform ablation studies to demonstrate the effectiveness of the graph router and the two coordination strategies used in GMoE. To verify the utility of each component, we individually remove the graph router, Poisson distinct loss (see in Eq. 7), and Normal balance loss (see in Eq. 9). The performance of the degraded variants, namely (-Graph), (-Normal), and (-Poisson), across four datasets is illustrated in Figure 2. We can see that all three components make contributions to the model’s performance. The degradation impact of removing the graph router (-Graph) is most obvious in ARC-Challenge and SIQA datasets in both model accuracy and stability aspects, indicating the importance of using the MoE graph to enhance expert collaborations. Moreover, the removal of the Poisson distribution loss (-Poisson) results in the most pronounced decrease in model performance on the BoolQ and OpenBookQA datasets, showing the essential role of maintaining the distinct capability of each expert. The normal distribution loss contributes to keeping the load balance of all experts and substantially improves the model performance.
5.4 Hyper-Parameter Analysis
@@
The performance of GMoE is affected by many hyper-parameters. We conduct analysis experiments to show the effects of hyper-parameters and present the results of ARC-Challenge and BoolQ datasets on Qwen2-7B in Fig. 3.
The number of experts.
@@
The number of experts varies from $\{4,8,12,16,32\}$. We can see that our model achieves the best performance with 8 experts. The increase in the number of experts may potentially complicate the collaboration process of experts and enlarge the imbalance problem.
Top-K experts.
@@
The number of experts activated by the router function with the Top-K largest assigned weights. The $K$ varies from $\{1,2,3,4,5\}$ in our experiments. We find our model performs the best with $K=2$.
Rank of LoRA.
@@
The hyper-parameter rank $r$ in LoRA controls the number of parameters that are updated in the fine-tuning process. The rank $r$ varies from $\{1,2,4,8,16,32\}$. Our model with $r=2$ achieves the best model performance which is less than the rank $r=4$ and $r=8$ required in MixLoRA in two datasets.
Edge density.
@@
The edge density controls the connection among expert nodes. The edge density is computed as the ratio of the number of edges in the MoE graph compared to the edges in the fully connected graph. The edge density $\beta$ varies from $\{10\%,20\%,40\%,60\%,80\%\}$. We find that with only $10\%$ connected edges, the collaboration information shared among experts facilitates our model to achieve the best results.
Initial edge connection.
In our experiments, the edges between all expert nodes are randomly constructed. To investigate whether the edge connections among experts in the initial graph affect the model’s performance, we repeat experiments under the same edge density $10\%$ with five random edge-connection settings. We can see that the initial connections among expert nodes have minimal influence on the performance of our model. Initially, all expert nodes are considered equal to obtain the input information. As training progresses, each expert node is automatically assigned corresponding weights based on the learning tasks to achieve optimal allocation effects.
5.4.1 Efficiency Analysis
As shown in Fig. 4, we quantify architectural efficiency through the number of Trainable parameters and Throughput (samples/second). Trainable parameters can be used as a critical metric for memory and storage costs, and Throughput can serve as a metric for computational speed that reflects the data processing rate during inference stages. These two metrics reflect the computational efficiency of the training and inference stages of our method. We can observe that our GMoE achieves the best model performance with the least number of trainable parameters. This phenomenon is primarily attributed to the much lower rank of LoRA required in our framework. The effective expert cooperation mechanism in our graph router reduces the number of parameters required by each expert, enabling our method to achieve both effectiveness and efficiency. In terms of model throughput, although our model does not have the highest throughput, it is comparable to other models while achieving the best model performance.